# Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies

### v2

@@ 3.3.3 Proposal 2: A Novel Estimand Based on Feasible Interventions
Unlike fixed interventions, which risk oversimplifying causal relationships by imposing uniform targets, the most feasible intervention approach offers a more nuanced and reasonable strategy. It acknowledges the inherent variability in real-world data and adapts the intervention accordingly, leading to more robust, reliable, and meaningful causal inferences that better reflect both the data and the practical realities of clinical and observational settings.
3.3.4 Data-Adaptive Trimming Causal Dose–Response Curve
In causal estimation, trimming refers to the practice of excluding observations with insufficient treatment support in order to stabilize estimation and mitigate the impact of positivity violations [7]. In its classical form, trimming relies on a fixed threshold $c>0$ applied to the conditional treatment density, yielding
$m^{\text{trimming}}(a):a\mapsto\mathbb{E}\!\left(Y^{a}\,\big|\,f_{A|\mathbf{L}}(a\mid\mathbf{L})>c\right),$ (17)
where $f_{A|\mathbf{L}}$ is the conditional density of $A$ given $\mathbf{L}$.
Estimand 4 (Trimming) In our framework, trimming is expressed in a more flexible, data-adaptive way using the empirical feasible set $\mathcal{A}_{\alpha}(\mathbf{L})$ introduced in Section 2.2. Specifically, we define $m^{\text{trimming}}(a):a\mapsto\mathbb{E}\!\left(Y^{a}\,\big|\,a\in\mathcal{A}_{\alpha}(\mathbf{L})\right),$ (18) that is, the mean counterfactual outcome restricted to those units for whom the intervention value $a$ lies within the empirically supported region given their covariates. This estimand can be identified by $\int\mathbb{E}(Y\mid A=a,\mathbf{L})\;t(a,\mathbf{L})\,dP_{\mathbf{L}}(\mathbf{l}),$ (19) where the trimming weight is $t(a,\mathbf{l})=\begin{cases}\dfrac{1}{\tau},&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{L}),\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (20) with $\tau=P_{0}\!\big(a\in\mathcal{A}_{\alpha}(\mathbf{L})\big)$ denoting the probability that $a$ falls within the empirical feasible region.
Comparison: This data-adaptive trimming estimand improves upon classical trimming by replacing an arbitrary, fixed threshold with a threshold derived from the observed distribution of the data. In doing so, it ensures that only well-supported units contribute to estimation, while retaining interpretability by conditioning on the feasible region. However, like classical trimming, it targets a restricted population rather than the full sample, which may reduce generalizability. In contrast, the feasible CDRC (Estimand 3) reassigns unsupported interventions rather than excluding them, preserving the entire study population while still avoiding extrapolation.
@@ 3.4 Estimation
3.4.1 Causal Estimation
To estimate the standard CDRC in Equation (6), a parametric g-formula can be readily applied for single time-point interventions. Substitution (plug-in) estimators can also be constructed to estimate the data-adaptive estimands introduced in Equation (15) and (18), with the full procedure detailed in Algorithm 1.
Notes:
@@ 3.4.1 Causal Estimation
    The estimated conditional treatment density $\hat{f}(a\mid\mathbf{l})$ (Step 2) is used only to define the data-adaptive estimands (feasible and trimming) and does not enter into the estimation of the outcome model.
@@
    Valid statistical inference using bootstrap intervals (Step 6) requires, under standard regularity, estimators that are $\sqrt{n}$-consistent and asymptotically linear with a well-defined influence function[3, 34, 35, 23]. These conditions are satisfied, for example, by correctly specified parametric outcome models, and can also be met by certain flexible estimators such as the Highly Adaptive Lasso (HAL) [2], which under mild conditions yields plug-in estimators that are $\sqrt{n}$-consistent and asymptotically linear in pathwise differentiable settings [6, 33]. The CDRC, however, is not pathwise differentiable as a full curve, so bootstrap validity cannot be guaranteed in general. At best, bootstrap may provide pointwise inference at fixed treatment values under correct parametric specification. In contrast, HAL plug-in estimators of the CDRC are pointwise asymptotically normal under undersmoothing, which enables valid delta-method inference based on the influence curve of the HAL coefficients [32]. This approach is analogous to the empirical sandwich variance estimator recently proposed for ICE g-computation by Zivich et al. [39].
@@
    When the target estimand is defined in a data-adaptive manner—such as via estimated highest density regions—standard independence assumptions are violated. As emphasized by Hubbard et al.[19], the intervention assigned to a given unit depends on functions of the data estimated from other units, introducing dependence between the observed data and the target parameter. This dependence may affect the validity of standard inferential procedures such as the bootstrap.
@@
    A proposed remedy is sample splitting or cross-validation, whereby the data-adaptive rule is defined on one subset of the data and the parameter is estimated on another. This restores approximate independence between the estimand and the estimation process. However, such approaches introduce practical complexity: they yield multiple, split-specific estimands, and averaging across them may reduce interpretability.
@@
    In this work, we choose to use the full sample to define and estimate the data-adaptive estimands. This preserves a single, coherent target parameter that reflects the region of intervention support within the observed data. While this choice introduces some additional dependence between the data and the estimand, we view it as a pragmatic balance between interpretability and statistical rigor.
  6. 6)
    Our simulation study (Appendix D) includes an empirical evaluation of bootstrap coverage, both with and without sample splitting. In the settings considered, bootstrap intervals achieve coverage close to nominal levels in well-supported regions. Thus, the practical benefit in interpretability outweighs the modest loss of conservativeness in inference for many applied scenarios.
@@ 3.4.2 Conditional Density Estimation
A key step in this process involves estimating the conditional density of the intervention given the confounders. Accurate conditional density estimation is essential for characterizing the regions in which the oracle intervention is supported by the data. Several approaches can be employed, including parametric models, nonparametric kernel conditional density estimators [12], and more flexible methods such as the Highly Adaptive Lasso–based density estimator (Haldensify) [13, 16, 14]. Motivated by Haldensify, in this paper we adopt a hazard-binning–based conditional density estimator, described in Appendix C, which is closely related to the approach of Díaz et al.[10]. This class of estimators can flexibly approximate conditional densities even when the relationship between confounders and interventions is complex or highly nonlinear.
Algorithm 1 Estimation Algorithm for $m^{\text{standard}},m^{\text{feasible}},m^{\text{trimming}}$, and $\tau$
1:Set-Up:
  • Let $n$ denote the number of subjects.
  • Define a grid of $m$ intervention values: $\overline{\mathbf{a}}=(a_{1},\cdots,a_{m})^{\mathrm{T}}$.
  • Choose a support level $\alpha$ to define highest density regions (HDRs).
2:Estimate Highest Density Regions:
  • (a)
    Conditional Density Estimation: Estimate the conditional treatment density $\hat{f}(a\mid\mathbf{l})$ using the observed data, and evaluate $\hat{f}(a_{i}\mid\mathbf{l}_{j})$ for all $i=1,\cdots,m$ and $j=1,\cdots,n$.
  • (b)
    Density Thresholds: For each subject $j$, determine $\hat{f}_{\alpha}^{j}$, the threshold such that the estimated HDR for subject $j$ contains probability mass $\alpha$.
3:Compute the Non-Overlap Ratio:
$\hat{\tau}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}I\big(\hat{f}(a_{i}\mid\mathbf{l}_{j})<\hat{f}_{\alpha}^{j}\big)$
4:Estimate the Outcome Model:
  • Estimate the conditional outcome model $\hat{E}(Y\mid A,\mathbf{L})$, for example using parametric regression.
  • Note: For plug-in estimation and bootstrap inference, we recommend outcome models with established theoretical guarantees (e.g., parametric models). Flexible methods such as HAL can also be considered.
5:Causal Estimation via Plug-in Estimators:
  • (a)
    Standard Estimand:
    $\hat{m}^{\mathrm{standard}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y\mid A=a_{i},\mathbf{l}_{j})$
  • (b)
    Feasible Estimand:
    • For each $a_{i}$ and subject $j$: if $\hat{f}(a_{i}\mid\mathbf{l}_{j})<\hat{f}_{\alpha}^{j}$, set $\hat{\mathop{}\!\mathrm{d}}(a_{i},\mathbf{l}_{j})$ to the nearest supported value within the estimated HDR; otherwise, set $\hat{\mathop{}\!\mathrm{d}}(a_{i},\mathbf{l}_{j})=a_{i}$.
    • Calculate
      $\hat{m}^{\mathrm{feasible}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}\big(Y\mid A=\hat{\mathop{}\!\mathrm{d}}(a_{i},\mathbf{l}_{j}),\mathbf{l}_{j}\big)$
  • (c)
    Trimming Estimand:
    • Estimate the trimming function:
      $\hat{t}(a_{i},\mathbf{l}_{j})=\begin{cases}\dfrac{1}{\hat{\tau}(a_{i})},&\text{if }\hat{f}(a_{i}\mid\mathbf{l}_{j})\geq\hat{f}_{\alpha}^{j},\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
    • Calculate
      $\hat{m}^{\mathrm{trimming}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y\mid A=a_{i},\mathbf{l}_{j})\cdot\hat{t}(a_{i},\mathbf{l}_{j})$
6:Construct Compatibility Intervals (Optional):
  • If using parametric or $\sqrt{n}$-consistent outcome models, and assuming standard regularity conditions, use the nonparametric bootstrap to construct approximate compatibility intervals for $\hat{m}^{\text{standard}}$, $\hat{m}^{\text{feasible}}$, and $\hat{m}^{\text{trimming}}$.

### v3

@@ 4.3 Novel Estimands to Address Positivity Violations: Feasible Dose–Response Curve (FDRC)
Unlike static interventions, which risk oversimplifying causal relationships by imposing uniform targets, the most feasible intervention approach offers a more nuanced and reasonable strategy. It acknowledges the inherent variability in real-world data and adapts the intervention accordingly, leading to more robust, reliable, and meaningful causal inferences that better reflect both the data and the practical realities of clinical and observational settings.
4.4 Trimmed dose–response curve
Estimand 4 (Trimmed Dose–Response Curve) Fix $\alpha\in(0,1)$. For each target level $a\in\mathcal{A}$, define $m_{0}^{\mathrm{trim}}(a):=\mathbb{E}_{0}\!\left[Y^{a}\mid a\in\mathcal{A}_{\alpha}(\mathbf{L};P_{0})\right],\qquad a\in\mathcal{A},$ (13) whenever $P_{0}\!\big(a\in\mathcal{A}_{\alpha}(\mathbf{L};P_{0})\big)>0$, equivalently $\tau_{0,\alpha}(a)<1$, where $\tau_{0,\alpha}$ the non-overlap ratio is defined in (5). Identification assumptions. For each $a\in\mathcal{A}$, assume: 1) Consistency: $Y=Y^{a}$ when $A=a$. 2) Conditional exchangeability: $Y^{a}\perp A\mid\mathbf{L}$. 3) Well-defined retained stratum: $P_{0}\!\big(a\in\mathcal{A}_{\alpha}(\mathbf{L};P_{0})\big)>0$ (equivalently $\tau_{0,\alpha}(a)<1$). Under 1)–3), this estimand is identified by $m_{0}^{\mathrm{trim}}(a)=\frac{\int\mathbb{E}_{0}\!\left(Y\mid A=a,\mathbf{L}=\mathbf{l}\right)\mathbbm{1}\!\{a\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})\}\,dP_{0,\mathbf{L}}(\mathbf{l})}{P_{0}\!\big(a\in\mathcal{A}_{\alpha}(\mathbf{L};P_{0})\big)}.$ Equivalently, $m_{0}^{\mathrm{trim}}(a)=\int\mathbb{E}_{0}\!\left(Y\mid A=a,\mathbf{L}=\mathbf{l}\right)\,t_{0,\alpha}(a,\mathbf{l})\,dP_{0,\mathbf{L}}(\mathbf{l}),$ where $t_{0,\alpha}(a,\mathbf{l})=\frac{\mathbbm{1}\!\{a\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})\}}{1-\tau_{0,\alpha}(a)}.$ Similar to FDRC, identification does not require positivity at level $a$ across all covariate strata; instead, by conditioning on $a\in\mathcal{A}_{\alpha}(\mathbf{L};P_{0})$, the estimand targets strata where $a$ is supported.
Comparison: Relative to conventional trimming estimands, this formulation replaces the fixed density threshold $c$ with a support restriction defined by the conditional HDR at level $\alpha$. This makes the notion of “sufficient support” explicit on the treatment scale through membership in $\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$, rather than through an arbitrary cutoff on $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})$.
As with any trimming estimand, the target parameter is conditional on a covariate-dependent retained set—here, strata with $a\in\mathcal{A}_{\alpha}(\mathbf{L};P_{0})$. This can improve statistical stability by reducing extrapolation to weak-support regions, but it also changes the causal target and may limit transportability to the full population.
@@ 4.5 Estimation
4.5.1 Estimation Algorithm
To estimate the standard CDRC in (6), a parametric g-formula can be applied directly for single time-point interventions. Substitution (plug-in) estimators can likewise be constructed for the population-level estimands in (12) and (13); the full procedure is given in Algorithm 1.
The complete CDRC and FDRC estimation workflow is implemented in the CICI package [31, 32], and illustrated in the Software Implementation section (Section 8). The software follows the same steps as Algorithm 1.
Algorithm 1 Plug-in estimation of $\hat{m}_{n}^{\mathrm{standard}}$, $\hat{m}_{n,\alpha}^{\mathrm{feasible}}$, $\hat{m}_{n,\alpha}^{\mathrm{trim}}$, and $\hat{\tau}_{n,\alpha}$ at target values
1:Inputs: $\{(\mathbf{L}_{j},A_{j},Y_{j})\}_{j=1}^{n}$; target grid $\overline{\mathbf{a}}^{\,\mathrm{tar}}=(a^{\mathrm{tar}}_{1},\ldots,a^{\mathrm{tar}}_{m_{\mathrm{tar}}})^{\top}\subset\mathcal{A}$; HDR grid $\overline{\mathbf{a}}^{\,\mathrm{hdr}}=(a^{\mathrm{hdr}}_{1},\ldots,a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}})^{\top}\subset\mathcal{A}$, assumed strictly increasing. Let $(b_{k})_{k=1}^{m_{\mathrm{hdr}}+1}$ be the induced cell boundaries defined by
$b_{1}:=a^{\mathrm{hdr}}_{1},\qquad b_{k}:=\frac{a^{\mathrm{hdr}}_{k-1}+a^{\mathrm{hdr}}_{k}}{2}\ \ (k=2,\dots,m_{\mathrm{hdr}}),\qquad b_{m_{\mathrm{hdr}}+1}:=a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}},$
and let $\{I_{k}\}_{k=1}^{m_{\mathrm{hdr}}}$ be the associated partition of $a^{\mathrm{hdr}}_{1},a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}}]$ given by
$I_{k}:=[b_{k},b_{k+1})\ (k=1,\dots,m_{\mathrm{hdr}}-1),\qquad I_{m_{\mathrm{hdr}}}:=[b_{m_{\mathrm{hdr}}},b_{m_{\mathrm{hdr}}+1}],$
with widths $w_{k}:=\lambda(I_{k})=b_{k+1}-b_{k}>0$. Support level $\alpha\in(0,1)$. Fix a deterministic tie-breaking rule for projections $\Pi_{T}(\cdot)$.
2:Estimate treatment density. Fit $\hat{f}_{n}(a\mid\mathbf{l})$ (see Section 4.5.2), and evaluate: $\hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{L}_{j})$ for all $k,j$, and $\hat{f}_{n}(a^{\mathrm{tar}}_{i}\mid\mathbf{L}_{j})$ for all $i,j$.
3:Construct subject-specific HDR cutoffs and supported sets. For each $j=1,\ldots,n$:
  1. (a)
    Compute $\hat{p}_{n,k}(\mathbf{L}_{j})$ by (21), using $w_{k}=\lambda(I_{k})$.
  2. (b)
    Compute $\hat{f}_{n,\alpha}(\mathbf{L}_{j})$ by (22).
  3. (c)
    Define the estimated supported set on the HDR grid:
    $\hat{\mathcal{A}}^{\,\mathrm{hdr}}_{n,\alpha}(\mathbf{L}_{j})=\Big\{a^{\mathrm{hdr}}_{k}\in\mathcal{A}:\ \hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{L}_{j})\geq\hat{f}_{n,\alpha}(\mathbf{L}_{j}),\ k=1,\ldots,m_{\mathrm{hdr}}\Big\}.$
4:Estimate non-overlap ratio at target values. For each $i=1,\ldots,m_{\mathrm{tar}}$,
$\hat{\tau}_{n,\alpha}(a^{\mathrm{tar}}_{i})=\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}\!\left\{\hat{f}_{n}(a^{\mathrm{tar}}_{i}\mid\mathbf{L}_{j})<\hat{f}_{n,\alpha}(\mathbf{L}_{j})\right\}.$
5:Estimate outcome regression. Fit $\hat{\mu}_{n}(a,\mathbf{l})\approx\mathbb{E}(Y\mid A=a,\mathbf{L}=\mathbf{l})$.
6:Compute plug-in curves at each target $a^{\mathrm{tar}}_{i}$. For each $i=1,\ldots,m_{\mathrm{tar}}$:
  1. (a)
    Standard CDRC:
    $\hat{m}_{n}^{\mathrm{standard}}(a^{\mathrm{tar}}_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{\mu}_{n}(a^{\mathrm{tar}}_{i},\mathbf{L}_{j}).$
  2. (b)
    Feasible FDRC:
    $\hat{d}_{n,\alpha}(a^{\mathrm{tar}}_{i},\mathbf{L}_{j})=\begin{cases}a^{\mathrm{tar}}_{i},&\hat{f}_{n}(a^{\mathrm{tar}}_{i}\mid\mathbf{L}_{j})\geq\hat{f}_{n,\alpha}(\mathbf{L}_{j}),\\ \Pi_{\hat{\mathcal{A}}^{\,\mathrm{hdr}}_{n,\alpha}(\mathbf{L}_{j})}(a^{\mathrm{tar}}_{i}),&\hat{f}_{n}(a^{\mathrm{tar}}_{i}\mid\mathbf{L}_{j})<\hat{f}_{n,\alpha}(\mathbf{L}_{j}),\end{cases}$
    $\hat{m}_{n,\alpha}^{\mathrm{feasible}}(a^{\mathrm{tar}}_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{\mu}_{n}\!\big(\hat{d}_{n,\alpha}(a^{\mathrm{tar}}_{i},\mathbf{L}_{j}),\mathbf{L}_{j}\big).$
  3. (c)
    Trimmed curve:
    $\hat{t}_{n,\alpha}(a^{\mathrm{tar}}_{i},\mathbf{L}_{j})=\frac{\mathbbm{1}\!\left\{\hat{f}_{n}(a^{\mathrm{tar}}_{i}\mid\mathbf{L}_{j})\geq\hat{f}_{n,\alpha}(\mathbf{L}_{j})\right\}}{1-\hat{\tau}_{n,\alpha}(a^{\mathrm{tar}}_{i})},$
    $\hat{m}_{n,\alpha}^{\mathrm{trim}}(a^{\mathrm{tar}}_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{\mu}_{n}(a^{\mathrm{tar}}_{i},\mathbf{L}_{j})\,\hat{t}_{n,\alpha}(a^{\mathrm{tar}}_{i},\mathbf{L}_{j}).$
7:(Optional) Uncertainty quantification. Use the nonparametric bootstrap for approximate compatibility intervals.
Remarks:
@@ 4.5.1 Estimation Algorithm
    The estimated conditional treatment density $\hat{f}_{n}(a\mid\mathbf{l})$ (Algorithm 1, Step 2) is used only to construct subject-specific cutoffs $\hat{f}_{n,\alpha}(\mathbf{l}_{j})$ and hence estimated supported sets
    $\hat{\mathcal{A}}^{\,\mathrm{hdr}}_{n,\alpha}(\mathbf{l}_{j})=\left\{a^{\mathrm{hdr}}_{k}\in\mathcal{A}:\hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{l}_{j})\geq\hat{f}_{n,\alpha}(\mathbf{l}_{j})\right\},$
    as well as derived quantities $\hat{d}_{n,\alpha}(a,\mathbf{l}_{j})$, $\hat{t}_{n,\alpha}(a,\mathbf{l}_{j})$, and $\mathbbm{1}\{a\in\hat{\mathcal{A}}^{\,\mathrm{hdr}}_{n,\alpha}(\mathbf{l}_{j})\}$ (Steps 6b–6c). It is not included as a covariate in the outcome regression $\hat{\mu}_{n}(a,\mathbf{l})$, except through these constructions.
@@
    The HDR cutoff computation used in Step 3 is the grid-based sample analogue of the population quantile characterization in Appendix C. Specifically, the population cutoff $f_{\alpha}(\mathbf{l})$ is defined as the $(1-\alpha)$-quantile of the density-value variable $Z_{\mathbf{l}}=f_{0}(A\mid\mathbf{l})$ in (19), and the corresponding HDR is the closed upper level set in (20). Algorithm 1 implements this via the weighted discrete approximation $\hat{p}_{n,k}(\mathbf{l})$ in (21), the estimated quantile $\hat{f}_{n,\alpha}(\mathbf{l})$ in (22), and the estimated closed upper level set $\hat{\mathcal{A}}^{\,\mathrm{hdr}}_{n,\alpha}(\mathbf{l})$ in (23). The grid-cell widths $w_{k}$ ensure proper normalization when the HDR grid is not equally spaced.
@@
    A formal consistency result for the feasible estimator is given in Appendix D and  E . Under Assumptions (A1)–(A8) in Section D.3, Theorem E.1 shows that for each fixed $a\in\mathcal{A}$,
    $\hat{m}^{\mathrm{feasible}}_{n,\alpha}(a)\xrightarrow{P}m^{\mathrm{feasible}}_{0,\alpha}(a).$
    The argument combines (i) consistency of $\hat{f}_{n}$ and $\hat{f}_{n,\alpha}$, which stabilizes estimated support membership (Lemma D.4.1); (ii) convergence of the estimated feasible rule $\hat{d}_{n,\alpha}(a,\mathbf{l})\to d_{0,\alpha}(a,\mathbf{l})$ (Lemma D.4.2); and (iii) consistency of $\hat{\mu}_{n}$, which controls plug-in regression error.
@@
    Step 2 accommodates multiple conditional-density estimators (e.g., parametric, binning, hazard-binning, HAL-based approaches). The hazard-binning construction and implementation details are provided in Section 8.2; the framework in Algorithm 1 is unchanged across these options.
@@
    Additional remarks on inference, bootstrap validity, first-stage nuisance estimation, and sample-splitting considerations are provided in Appendix F.
@@ 4.5.2 Conditional Density Estimation
A key step is estimation of the conditional treatment density $f_{A\mid\mathbf{L}}(a\mid\mathbf{l})$, which determines supported intervention regions and subject-specific feasibility cutoffs. Several estimators can be used, including parametric models, nonparametric kernel conditional density estimators [12], and flexible machine-learning approaches such as HAL-based density estimation (Haldensify) [14, 15, 13].
In our implementation, we allow multiple density-estimation options within the same plug-in framework. One option is a hazard-binning estimator motivated by Haldensify and related to the pooled-hazard approach of Díaz Muñoz and van der Laan [10]; its construction is described in Section 8.2. Regardless of the specific estimator used, the downstream steps in Algorithm 1 (HDR construction, feasible mapping, trimming weights, and plug-in evaluation) remain the same.
Remark 4.1 (Grid-based versus closed-form HDR construction).
Some conditional density estimators are naturally implemented on a prespecified grid in $a$. In particular, grid-based methods such as Haldensify and the hazard-binning estimator (Section 8.2) return $\hat{f}_{n}(a\mid\mathbf{l})$ through evaluations on a finite grid, so we construct the HDR and the subject-specific cutoff using the HDR grid $\overline{\mathbf{a}}^{\,\mathrm{hdr}}$ together with its induced cell widths $\{w_{k}\}$ (Appendix C.2).
In contrast, for fully parametric conditional densities (e.g., Gaussian regression), the HDR can in principle be characterized directly as an upper level set of $\hat{f}_{n}(\cdot\mid\mathbf{l})$ without discretization. Nevertheless, even in the parametric case, the same grid-based construction is often convenient and numerically stable, and it matches the pragmatic implementation used in Algorithm 1.